I have two questions that I missed on a test. Will someone please help me.
1. The sum of a number and 17 more than twice the same number is 101. Find the number. I came up with 67.
2. Write an equation in slope-intercept form for the line that passes through (3,5) and (-2,1).
I must show my work to get credit for these two problems.
Thanks

Answer by Alan3354(30993)

I have two questions that I missed on a test. Will someone please help me.
1. The sum of a number and 17 more than twice the same number is 101. Find the number. I came up with 67.
-------------------
n = number
n + (2n+17) = 101
3n + 17 = 101
3n = 84
n = 28
How did you get 67?
-----------------
2. Write an equation in slope-intercept form for the line that passes through (3,5) and (-2,1).
First, the long way.
Find the slope, m, = (diff in y)/(diff in x)
m = (5-1)/(3-(-2))
m = 4/5
Then use either point
y-y1 = m*(x-x1)
y-5 = (4/5)*(x-3)
y-5 = (4/5)x - 12/5
y = (4/5)x + 13/5 (in slope-intercept form)
----------------
Easier method:
|x y +1|
|3 5 +1| = 0
|-2 1 1|
Note: the + signs aid in alignment
x*(5-1) - y*(3 -(-2)) + (3-(-10)) = 0
4x -5y +13 = 0
5y = 4x + 13
y = (4/5)x + 13/5
--------------
I must show my work to get credit for these two problems.